How many moles of MgSiO3 are in 207 g of the compound?
1
Expert's answer
2015-10-29T07:24:17-0400
To get the number of moles we should divide the grams given by the substance's molar mass. The molar mass (M) of MgSiO3 is 24.305 + 28.086 + 3x15.999 = 100.389 g/mol.
The amount of substance is n(MgSiO3) = m(MgSiO3)/M(MgSiO3) = 207/100.389 = 2.062 mol

Answer: 2.062 mol